OpenStudy (thatcreepyboy):

My math skills are terrible Box A has a volume of 12 cubic meters. Box B is similar to box A. To create Box B, Box A's dimensions were multiplied by five. What is the volume of Box B? 60 m3 300 m3 1,500 m3 7,500 m3

OpenStudy (thatcreepyboy):

MEDAL AND FAN

OpenStudy (thatcreepyboy):

EXTRA MEDAL AND FAN PEOPLE

OpenStudy (will.h):

Well I think Since the dimensions of A were multiplied by 5 to create B then you will have to multiply the volume of Box A by 5 in order to find the volume of Box B Box B volume = 5 * box A volume Box B volume = 5 * 12 Box B volume = 60

OpenStudy (thatcreepyboy):

okay THANK YOU